OpenStudy (anonymous):

Need help with a mathematical induction problem. 1+2+2^2+2^3...+2^(n-1) =2^(n) -1. I keep getting down to 4^(k-1) but It doesn't look like the answer. I plugged in (k+1) into 2^(n) -1 to know what I'm supposed to end up with to prove this formula works for the k+1 term.

OpenStudy (tkhunny):

Does it start? 1 + 2 + 4 = 7 2^3 - 1 = 8 - 1 = 7 Does one imply the next? 1+2+2^2+2^3...+2^(n-1) + 2^n 2^(n) -1 + 2^n = 2(2^n) - 1 = 2^(n+1) - 1 And we appear to be done. Why do you have a four (4) in that base?

OpenStudy (anonymous):

I don't understand how you simplified from the last step. 2^(n) -1 + 2^n = 2(2^n) - 1 = 2^(n+1) - 1 I added the two 2^(k) and then put -1 to get 4^(K) -1

OpenStudy (tkhunny):

?? \(2^{n} + 2^{n} = 2^{n}(1+1) = 2\cdot 2^{n} = 2^{n+1}\) \(House + House = 2*House \ne Duplex\) Why would you add or multiply the bases?